OpenStudy (anonymous):

Solve the system algebraically. Check your work. 5x + 2y = 10 3x + 2y = 6 Make sure there are NO SPACES in your answer. Be sure to include a comma.

OpenStudy (anonymous):

I don't get this question at all...

OpenStudy (anonymous):

5x + 2y = 10 3x + 2y = 6 ---->(-1) -------------- 5x + 2y = 10 -3x - 2y = -6 (result of multiplying by -1) -------------add 2x = 4 -- divide by 2 x = 2 Sub in 2 for x in either equation to find y 5x + 2y = 10 5(2) + 2y = 10 10 + 2y = 10 2y = 10 - 10 2y = 0 y = 0 check.. 3x + 2y = 6 3(2) + 2(0) = 6 6 + 0 = 6 6 = 6 (correct) solution : (2,0)

OpenStudy (anonymous):

Thank you very much! you explained it very well

OpenStudy (anonymous):

your welcome. This can also be solved by using the substitution method. The reason I didn't use that way was because I do not like dealing with fractions

OpenStudy (whpalmer4):

You can do substitution on this without doing fractions. \[5x + 2y = 10\]\[ 3x + 2y = 6 \]We have \(2y\) in both equations, so solve one for \(2y\) and substitute: \[5x+2y=10\]\[2y=10-5x\]\[3x+(10-5x) = 6\]\[-2x+10=6\]\[-2x=-4\]\[x=-2\]etc.

OpenStudy (whpalmer4):

sorry, obviously that should be \[x = 2\]!